Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 mph and train B is traveling at 50 mph. Train A passes a station at 9:10 AM. If train B passes the same station at 9:22 am, at what time will train B catch up to Train A?
Answer by mananth(16949)

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 mph and train B is traveling at 50 mph. Train A passes a station at 9:10 AM. If train B passes the same station at 9:22 am, at what time will train B catch up to Train A?


train A is at station at 9.10 am
Train B is at the station at 9.22 am
in 12 minutes train A has traveled 8 miles ( 1/5 hour *40 mph)
Let train B travel x miles from the station when it catches up with train A
time taken by A to travel the distance x-8 miles = x-8 / 40
time taken by B to travel x miles is x/50
Both these times are same x/50 = x-8 /40
40x= 50( x-8)
40x = 50x-400
40x-50x=-400
-10x= -400
x= -400 / -10
x= 40 miles.
40 / 50 is the time taken by B to travel from station when it catches up with A
= 4/5 hours = 48 minutes
9.22 am + 48minutes = 10.10 am
At 10.10 am train B will catch up with train A